OpenStudy (smurfy14):

Solve for θ_subscript2

OpenStudy (smurfy14):

\[n _{1}\sin \theta _{1}=n _{2}\sin \theta _{2}\]

hero (hero):

Oops

hero (hero):

\[\sin^{-1}(\frac{n_1 \sin \theta_1}{n_2}) = \theta_2\]

OpenStudy (smurfy14):

again, can u show me the steps

OpenStudy (anonymous):

Divide through by n2, then take the inverse sine of both sides.

hero (hero):

First, I divided both sides by n_2. Next I took the inverse sin of both sides. (Not sure if that's the correct terminology, but...any who

OpenStudy (anonymous):

That's why hero is a superhero

hero (hero):

I'm just an amateur. Nobody special. But I aspire to be a superhero.

hero (hero):

Mandolino, I'll give you a medal for good intention